OpenStudy (anonymous):

Which is the solution set for –7 ≤ 4x + 1 ≤ 5? A. x ≤ –2 or x ≥ 1 B. x ≥ –2 and x ≤ 1 C. x ≥ and x ≤ D. x ≥ or x ≤

OpenStudy (abb0t):

What is the last two? You're missing information.

OpenStudy (anonymous):

sry here C. x ≥- 3/2 and x ≤3/2

OpenStudy (anonymous):

-7 <= 4x + 1 <= 5 --- subtract 1 -7 - 1 <= 4x + 1 - 1 <= 5 - 1 -8 <= 4x <= 4 --- divide by 4 -8/4 <= 4x/4 <= 4/4 -2 <= x <= 1

OpenStudy (anonymous):

D. x ≥-3/2 or x ≤ 3/2

OpenStudy (anonymous):

answer is B

OpenStudy (anonymous):

Thanks! i hate asking for help.

OpenStudy (anonymous):

no worries....everybody needs help sometimes :)

OpenStudy (anonymous):

yea. thanks!